# Exact Consistency Under Partial Views: Graph Colorability, Capacity, and Equality in Multi-Location Encodings

### 16 Mar 2026 v3

@@ IV Unified Finite Converse
Sufficiency. If $\phi$ is injective, define a decoder on the image of $\phi$ by $D(y,t)=\phi^{-1}(y,t)$, with arbitrary extension off the image. Injectivity makes $\phi^{-1}$ single-valued on the image, and for every latent state $x$ we then have $D(Y(x),T(x))=x$. Thus exact zero-error recovery is possible.  $\blacksquare$
Proposition IV.2 (Confusability formulation).
Fix a deterministic observation transcript and an $L$-bit side-information tag. If $K$ latent states induce the same observation transcript, then exact zero-error decoding on that ambiguity class requires $K$ distinct tag outcomes. Equivalently, a $K$-way ambiguity class forms a confusability clique whose size cannot exceed the available tag alphabet.
(L: OBS5)
@@
Apply Theorem IV.1 on an ambiguity class on which the observation is constant. Then injectivity of $(Y,T)$ reduces to injectivity of the tag coordinate alone on that class. Since an $L$-bit tag provides at most $2^{L}$ outcomes, the clique size cannot exceed $2^{L}$.  $\blacksquare$
Global observation-tag budget.
If the observation alphabet has size $O$ and the auxiliary tag alphabet has size $T$, then exact zero-error recovery on a latent alphabet of size $K$ requires
$K\leq O\,T.$
(L: OBS4)
Proof.
Let $\phi(x)=(Y(x),T(x))$. By Theorem IV.1, exact zero-error recovery requires $\phi$ to be injective on the latent alphabet under consideration. The codomain of $\phi$ is the finite product alphabet $\mathcal{Y}\times\mathcal{T}$, which has cardinality $|\mathcal{Y}\times\mathcal{T}|=OT$. An injective map from a set of size $K$ into a set of size $OT$ can exist only if $K\leq OT$. This is exactly the claimed global budget bound.  $\blacksquare$
Fiber-level injectivity and cardinality.
Fix an observation value $y$. On the observation fiber
$\mathcal{F}_{y}=\{x:Y(x)=y\},$
exact zero-error recovery forces the tag map to be injective. Consequently,
$|\mathcal{F}_{y}|\leq|\mathcal{T}|.$
(L: OBS3)
Proof.
Fix $y$ and restrict attention to the fiber $\mathcal{F}_{y}=\{x:Y(x)=y\}$. On this set the observation coordinate is constant, so for $x,x^{\prime}\in\mathcal{F}_{y}$ one has
$(Y(x),T(x))=(Y(x^{\prime}),T(x^{\prime}))\iff T(x)=T(x^{\prime}).$
Hence Theorem IV.1 implies that exact recovery on $\mathcal{F}_{y}$ is possible only if the tag map $x\mapsto T(x)$ is injective on that fiber. Since the image of an injective map cannot be larger than its codomain, one obtains $|\mathcal{F}_{y}|\leq|\mathcal{T}|$.  $\blacksquare$
Proposition IV.3 (Finite counting converse).
Let $F$ range over $K$ possible latent values. Suppose exact recovery is attempted from a fixed observation transcript together with an $L$-bit side-information tag. Then
$K\leq 2^{L}.$
Equivalently, exact zero-error recovery of $K$ ambiguous states requires at least $\log_{2}K$ bits of side information.
(L: CIA3)
Proof.
@@
This is Theorem IV.2 specialized to a single $K$-way ambiguity class.  $\blacksquare$
The counting bound is the coarsest form of the obstruction. The next step is to weight the same argument by source mass. Exact recovery still enforces injectivity on successful confusability classes, but now the question is how much entropy can remain once the source is partitioned into success and failure branches.
Conditional-entropy formulation.
Let $X$ be uniform on a $K$-way ambiguity class and let $Y$ be the deterministic observation transcript. If $Y$ is constant on that class, then $H(X\mid Y)=\log_{2}K$. In particular, any tag-observation pair $(Y,T)$ that resolves the class exactly must satisfy
$H(X\mid Y,T)=0\qquad\text{and}\qquad H(T\mid Y)\geq\log_{2}K.$
(L: PRB47, PRB55)
Proof.
If $Y$ is constant on the $K$-point ambiguity class, then conditioning on $Y$ does not refine the distribution of $X$. Because $X$ is uniform on that class, this gives $H(X\mid Y)=\log_{2}K$. Exact recovery from $(Y,T)$ means that $X$ is a deterministic function of $(Y,T)$, so $H(X\mid Y,T)=0$. Using the chain rule,
$H(X\mid Y)=I(X;T\mid Y)+H(X\mid Y,T)=I(X;T\mid Y).$
Therefore $I(X;T\mid Y)=\log_{2}K$. Since conditional mutual information is bounded by conditional entropy, $I(X;T\mid Y)\leq H(T\mid Y)$, and the lower bound $H(T\mid Y)\geq\log_{2}K$ follows.  $\blacksquare$
Mass-weighted clique entropy bound.
Let $S$ be a success set for a zero-error decoder, and suppose $S$ forms a confusability clique under the observation transcript. If $p_{S}$ denotes the total probability mass of $S$, then the entropy carried by the successful states obeys
$\sum_{x\in S}p(x)\log_{2}\frac{1}{p(x)}\;\leq\;h_{2}(p_{S})+p_{S}\log_{2}|\mathcal{T}|,$
where $h_{2}$ is binary entropy and $|\mathcal{T}|$ is the tag alphabet size. (L: PRB55)
Proof.
Let $B$ be the indicator of the event $\{X\in S\}$. Then $\Pr[B=1]=p_{S}$, so $H(B)=h_{2}(p_{S})$. Decompose the source entropy carried by the successful states by conditioning on $B$:
$H(X)=H(B)+H(X\mid B).$
Restricting to the success branch $B=1$, exact recovery on the clique $S$ forces injectivity of the tag map on $S$, hence at most $|\mathcal{T}|$ successful states can be resolved. Therefore
$H(X\mid B=1)\leq\log_{2}|\mathcal{T}|.$
Multiplying by the success probability gives a contribution of at most $p_{S}\log_{2}|\mathcal{T}|$ from that branch. Combining this with the Bernoulli split term $h_{2}(p_{S})$ yields the displayed inequality for the entropy mass carried by the successful states.  $\blacksquare$
In particular, on any exact success set inside a confusability clique, the restricted observation-tag map is injective by Theorem IV.1, so the same entropy inequality applies directly to the successful branch rather than only to the whole clique.(L: OBS1, PRB55)
Decoder-output and gap formulations.
In the same deterministic finite model, the obstruction can also be expressed as output-entropy and finite-budget gap constraints: any deterministic decoder output $\hat{X}$ satisfies
$H(\hat{X})\leq H(Y,T)\leq H(X),$
and the observation-tag entropy and decoded-output entropy are bounded by their corresponding finite alphabet ceilings, so the deficits
[…]
are nonnegative and vanish only in the uniform saturation cases formalized in the supplement. (L: PRB73-74, PRB82-83, PRB90, PRB93-95)
Proof.
Because the decoder is deterministic, $\hat{X}=D(Y,T)$ is a deterministic function of the pair $(Y,T)$. The deterministic data-processing statement below therefore gives
$H(\hat{X})\leq H(Y,T).$
The pair $(Y,T)$ is itself a deterministic function of the latent source $X$, so another application of deterministic data processing yields $H(Y,T)\leq H(X)$. Since $(Y,T)$ takes values in the finite alphabet $\mathcal{Y}\times\mathcal{T}$, one also has
$H(Y,T)\leq\log_{2}|\mathcal{Y}\times\mathcal{T}|.$
[…]
Deterministic data processing.
Let $\kappa$ be any deterministic coarsening of the observation-tag pair $(Y,T)$. Then
$H(\kappa(Y,T))\leq H(Y,T).$
In particular, since any deterministic decoder output $\hat{X}$ is a coarsening of $(Y,T)$,
$H(\hat{X})\leq H(Y,T)\leq\log_{2}|\mathcal{Y}\times\mathcal{T}|.$
(L: PRB68, PRB81, PRB83)
Proof.
Write $Z=(Y,T)$. If $\kappa$ is deterministic, then $\kappa(Z)$ is a function of $Z$, so the conditional entropy $H(\kappa(Z)\mid Z)$ is zero. By the chain rule,
$H(Z)=H(\kappa(Z),Z)=H(\kappa(Z))+H(Z\mid\kappa(Z)),$
and since conditional entropy is nonnegative, $H(\kappa(Z))\leq H(Z)$. Taking $\kappa$ to be the decoder map gives $H(\hat{X})\leq H(Y,T)$. Finally, because $Z$ ranges over the finite alphabet $\mathcal{Y}\times\mathcal{T}$, one has the standard finite-alphabet ceiling $H(Y,T)\leq\log_{2}|\mathcal{Y}\times\mathcal{T}|$.  $\blacksquare$
This theorem is the structural reason the output and gap formulations are not independent embellishments of the counting converse. Once the observation is coarsened, every derived representation inherits the same obstruction through deterministic data processing rather than escaping it.(L: PRB68, PRB81, PRB83, PRB90)
Proposition IV.4 (Equivalence viewpoint).
@@
The confusability, counting, conditional-entropy, decoder-output, and finite-gap statements above are different normal forms of the same deterministic finite zero-error obstruction: a surviving $K$-way ambiguity class requires a budget of at least $\log_{2}K$ bits to be resolved exactly.
(L: OBS1, OBS5, CIA3, PRB47, PRB55, PRB68, PRB81, PRB83, PRB90)
Proof.
Pair injectivity is the structural core. The confusability and counting theorems express the injectivity requirement combinatorially; the conditional-entropy and weighted-entropy theorems express it in source-mass coordinates; the decoder-output and gap theorems express it through deterministic coarsening and finite alphabet ceilings; and the finite-error theorem is the relaxed version obtained by allowing a nonzero failure branch. Each theorem therefore measures the same failure of exact isolation in a different coordinate system.  $\blacksquare$
IV-A Finite-error extension
@@
The main body of the paper studies exact zero error. The same deterministic finite model also supports a budgeted finite-error extension, which should be read as a deterministic finite analogue of the classical Fano line of argument rather than as a separate asymptotic coding theorem [8, 9, 6].
Proposition IV.5 (Budgeted finite-error converse).
@@
V Graph Characterization of Partial Views
@@
V-A A Running Example: The Binary Square
@@
Suppose the architecture exposes two partial views:
  • View 1 reveals only coordinate $x_{1}$
  • View 2 reveals only coordinate $x_{2}$
@@
Two states are confusable when at least one view fails to separate them:
  • Through View 1: states $(0,0)$ and $(0,1)$ are confusable (both have $x_{1}=0$), and states $(1,0)$ and $(1,1)$ are confusable (both have $x_{1}=1$).
  • Through View 2: states $(0,0)$ and $(1,0)$ are confusable (both have $x_{2}=0$), and states $(0,1)$ and $(1,1)$ are confusable (both have $x_{2}=1$).
The confusability graph is therefore a 4-cycle:
@@
where opposite corners $(0,0)$ and $(1,1)$, as well as $(0,1)$ and $(1,0)$, are not adjacent. This is not a clique: the architecture can distinguish some pairs even though it cannot distinguish all pairs.
@@
Because the graph is 2-colorable (e.g., color by parity $c(x_{1},x_{2})=x_{1}\oplus x_{2}$), exact recovery is possible with a 2-ary auxiliary tag. However, the graph is not 1-colorable (it contains edges), so a 1-tag decoder must fail on at least one confusable pair. Under the uniform source, the best 1-tag exact success probability is $1/2$ (choose either diagonal as the success set).
What this example shows.
  1. 1.
    Partial views induce structured confusability, not uniform ambiguity.
  2. 2.
    The confusability graph captures exactly which state pairs the architecture cannot separate.
  3. 3.
    Graph-theoretic properties (chromatic number, independence number) determine exact recovery costs.
@@
Coding-theoretic interpretation. This 4-cycle structure has a direct interpretation in error-correcting codes. Coloring by parity $c(x_{1},x_{2})=x_{1}\oplus x_{2}$ is exactly a syndrome: the tag reveals whether the two bits match. With one syndrome bit ($T=2$), we can decode perfectly because each color class contains non-adjacent states. With no tag ($T=1$), we cannot distinguish adjacent vertices, exactly as in the classic binary symmetric channel without feedback. The confusability graph is precisely the confusion graph of a code with parity-check matrix H=[1 1]H=[1\;1]: adjacent vertices are those the code cannot separate.
@@
Figure 1 contrasts this structured failure with the clique-shaped failure that arises when no observation provides any discrimination. The theorems below generalize this phenomenon to arbitrary multi-fact partial-view systems.
V-B General Graph Characterization
This section turns the one-shot obstruction into the paper’s main structural object: the confusability graph induced by admissible partial views. The single-fact converse family treats ambiguity classes that are effectively clique-shaped under the observation transcript. One extension is to let the latent state be a tuple of facts and let each location reveal only a subset of coordinates. Failure then becomes structured: some latent pairs remain indistinguishable and others do not, so the resulting confusability graph need not be complete.
In this extension, ambiguity is now over full latent tuples rather than single fact values: two tuples are confusable when the allowed partial observations fail to separate them. The induced graph is therefore a failure map for the architecture.
The graph can be handled either explicitly or implicitly. If there are $d$ facts over an alphabet of size $q$, then the latent state space has cardinality $q^{d}$, so any explicit confusability graph materialization already starts from an exponentially large vertex set.(L: MFT113) A naive explicit construction ranges over at most $q^{2d}$ ordered state pairs.(L: MFT116-117) For that reason the appropriate algorithmic interface is implicit: the formalization packages adjacency as a Boolean confusability oracle on state pairs, proves that this oracle is equivalent to the confusability relation itself, and also packages an equivalent agreement-set oracle together with a linear bound on the total view-membership work needed by the direct scan implementation.(L: MFT114-115, MFT132-135) Under the standard explicit representation in which a latent tuple is stored as a length-$d$ coordinate array and each admissible view is stored as its coordinate subset, this one-shot oracle is therefore computable by direct inspection in deterministic time $O(d+\sum_{\ell}|V_{\ell}|)$, hence $O(Ld)$ for $L$ views: compute the agreement set of the two tuples once, then scan the views and test whether some view is contained in that agreement set. The theory below is therefore compatible with implicit graph access even when explicit graph materialization is impractical, while still leaving the one-shot adjacency test itself polynomial in the size of its explicit input.
@@ V-B General Graph Characterization
The deterministic restriction is also not fundamental at the graph-construction level. The Lean development formalizes a support-overlap extension in which each location may emit any observation from a finite support set, and two states are confusable when some admissible location has overlapping observation support on those states. The present deterministic partial-view model embeds into that broader framework as the singleton-support special case, so the graph construction itself already extends beyond exact deterministic views even though the paper develops only the deterministic architectural arc.(L: MFT110-112)
At the same time, the exact current coordinate-subset model is not graph-universal on the full tuple space, and the induced graph class can be characterized more sharply than mere vertex transitivity. The formalization first shows that equality at a view is equivalent to containment of that view inside the coordinate-agreement set of the two tuples, so confusability depends only on which coordinates agree, not on the actual symbols carried there.(L: MFT118-119, MFT122) From any view family one therefore obtains an upward-closed family of coordinate subsets,
$\mathcal{U}:=\{S\subseteq[F]:\exists\ell\ \text{with}\ V_{\ell}\subseteq S\},$
@@
such that two distinct tuples are adjacent if and only if their agreement set lies in $\mathcal{U}$. Conversely, every upward-closed family $\mathcal{U}$ arises from some coordinate-view architecture, so on the full labeled tuple space the realizable confusability relations are exactly the agreement-set relations cut out by upward-closed families.(L: MFT125-128) When the alphabet has at least two symbols, every proper coordinate subset occurs as the agreement set of some distinct pair of tuples, so this characterization is complete up to the irrelevant full-agreement case corresponding to identical tuples.(L: MFT129-131) Equivalently, coordinatewise alphabet permutations preserve the confusability relation and induce graph automorphisms; for any two latent tuples there exists such an automorphism carrying one to the other.(L: MFT120-124) Thus the exact model realizes a specific monotone agreement-set graph class rather than an arbitrary finite graph family. In particular, non-vertex-transitive graphs such as a three-vertex path cannot arise as full confusability graphs in this exact model. What remains open is a classification up to unlabeled graph isomorphism, and the corresponding classification for restricted realized-state subsets or for the stochastic support-overlap extension.
Theorem V.1 (Exact recovery as graph colorability).
@@
For a finite multi-fact latent tuple observed through a finite family of partial-coordinate views, exact zero-error recovery with a $T$-ary auxiliary tag exists if and only if the induced confusability graph is $T$-colorable.
(L: MFT3-4, MFT20)
@@
Necessity. Suppose exact recovery with $T$ tags is possible. If two latent tuples are adjacent in $G_{\mathrm{conf}}$, then by definition some allowed observation leaves them indistinguishable. Exact recovery therefore cannot assign them the same tag, because the joint observation-tag pair would then coincide on two distinct latent tuples and Theorem IV.1 would be violated. Hence the tag assignment is a proper $T$-coloring of $G_{\mathrm{conf}}$.
Sufficiency. Suppose $c$ is a proper $T$-coloring of $G_{\mathrm{conf}}$. Define the auxiliary tag of a latent tuple to be its color. If two latent tuples produce the same observation and the same color, then they are confusable and adjacent, contradicting proper colorability. Therefore the joint observation-color map is injective, and Theorem IV.1 yields an exact decoder.  $\blacksquare$
An edge of $G_{\mathrm{conf}}$ records a specific pair of latent states that the view architecture cannot separate exactly, and a proper coloring is exactly a disambiguation budget that resolves those failures. Exact recovery cost is determined by failure topology, not by the count of sources alone.
This statement is packaged through an explicit simple confusability graph object, and for $n$ repeated copies the full block confusability graph is identified with the $n$-fold strong power of the one-shot confusability graph. The zero-error criterion is therefore literally a graph-colorability theorem on an explicit power family rather than only an analogy.(L: MFT34-35)
Theorem V.2 (Success-set characterization).
@@
Restrict the argument of Theorem V.1 to the success set. Exactness is required only on $S$, so the coloring condition is required only on confusable pairs inside $S$.  $\blacksquare$
Theorem V.3 (Exact finite-error success budget).
@@
$M_{T}:=\max\{|S|:S\text{ induces a }T\text{-colorable confusability subgraph}\}.$
@@
Theorem V.2 identifies exact success sets with induced $T$-colorable subgraphs. Since the latent alphabet is finite, a maximum-cardinality such set exists. The mechanized characterization exposes this optimum directly. In the binary four-cycle witness, the optimum at $T=1$ is exactly two states, so the best uniform exact success probability is $2/4=1/2$.  $\blacksquare$
The same characterization is also available in weighted form. For any finite source weight $w$ on the latent alphabet, the maximum mass of a $T$-colorable induced subgraph is the exact finite success value under budget $T$. This is the weighted version of the same structural claim: under limited budget, the best exact decoder is the largest colorable safe subworld allowed by the induced failure graph.(L: MFT18-19)
Equivalently, this optimum can be packaged as an explicit exact finite rate-distortion value for the extended model at tag budget $T$: every exact success set has weighted mass at most this value, and some exact decoder attains it. Thus the extended model no longer has only a converse inequality; at the one-shot level it has an exact weighted success law rather than merely an upper bound.(L: MFT23-24)
Theorem V.4 (First non-clique witness).
@@
Theorem V.5 (Block-composition law).
@@
(L: MFT11-13, MFT16)
For later use, recall the strong product definition: if $G$ and $H$ are graphs, then $G\boxtimes H$ has vertex set $V(G)\times V(H)$, and $(u,v)$ is adjacent to $(u^{\prime},v^{\prime})$ exactly when either $u=u^{\prime}$ and $v\sim v^{\prime}$, or $u\sim u^{\prime}$ and $v=v^{\prime}$, or $u\sim u^{\prime}$ and $v\sim v^{\prime}$. The block-composition theorem is model-specific because the admissible product views generate exactly these three adjacency cases.
@@
This product law now has both directions in the formal development. The upper direction is the multiplicative coloring construction above. More sharply, once both location alphabets are nonempty, the product confusability graph is exactly the strong product of the two component confusability graphs. The graph object is therefore not just analogous to zero-error graph products; it is literally a strong-product construction for the induced failure map. (L: MFT25-26)
The lower direction is clique-based: if the first component contains a confusability clique of size $c_{1}$ and the second contains one of size $c_{2}$, then the product system contains a confusability clique of size $c_{1}c_{2}$, so any exact product decoder requires at least $c_{1}c_{2}$ tags. The same mechanism also propagates to finite-error success sets: the largest exactly decodable success set in the product system is at least the product of the largest exactly decodable success sets in the two components. At the one-shot zero-error coding level, the maximum independent-set code size also obeys the same product lower bound. Thus the extended model now has an honest strong-product graph law together with matching lower/upper budget consequences.(L: MFT17)
At this point the induced graph class has four explicit structural features. It contains non-clique graphs, as witnessed by the binary square. It contains the cluster-collapse subclass on which the later equality theorem is exact. It is closed under the block-composition operation, which becomes strong product on confusability graphs. And, by the agreement-set characterization above, it is exactly a monotone agreement-set class rather than an arbitrary graph family on the full tuple space. The iterated family below packages these features into a concrete infinite class.
@@
The four-state witness above is only the smallest member of a larger model-generated family. Let $W$ denote the binary square system of Theorem V.4. Composing $m$ independent copies of $W$ yields a system with $2m$ binary fact coordinates and $4^{m}$ latent states whose confusability graph is the $m$-fold strong power
$C_{4}^{\boxtimes m}.$
Because each copy of $W$ is exactly recoverable with two tags, repeated application of Theorem V.5 gives exact recovery on the $m$-fold product with $2^{m}$ tags. Conversely, each copy contains a confusability clique of size $2$, so the clique lower bound in the same product theorem yields a confusability clique of size $2^{m}$ in the product system. Hence every exact decoder for this family requires at least $2^{m}$ tags, and therefore the exact tag budget is
$T_{\mathrm{exact}}(W^{\otimes m})=2^{m}.$
This provides an infinite family of non-clique examples, generated directly by the model, on which the graph object, the exact budget, and the strong-product law all interact nontrivially rather than only in the single four-cycle case. At this point the paper has moved decisively beyond the trivial threshold story: above the unit-rate corner, not all failures are equally severe, because the architecture can induce a non-clique topology with a strictly more structured recovery law. (L: MFT11-13, MFT16-17, MFT29-30)
@@
VI Block Composition, Strong Powers, and Asymptotic Capacity

### 31 Mar 2026 v4

@@ III Unified Finite Converse
Sufficiency. If $\phi$ is injective, define a decoder on the image of $\phi$ by $D(y,t)=\phi^{-1}(y,t)$, with arbitrary extension off the image. Injectivity makes $\phi^{-1}$ single-valued on the image, and for every latent state $x$ we then have $D(Y(x),T(x))=x$. Thus zero-error recovery is possible.  $\blacksquare$
Theorem III.2footnotemark: footnotetext: : OBS5 (Confusability formulation).
Fix a deterministic observation transcript and an $L$-bit side-information tag. If $K$ latent states induce the same observation transcript, then zero-error decoding on that ambiguity class requires $K$ distinct tag outcomes. Equivalently, a $K$-way ambiguity class forms a confusability clique whose size cannot exceed the available tag alphabet.
@@
Apply Theorem III.1 on an ambiguity class on which the observation is constant. Then injectivity of $(Y,T)$ reduces to injectivity of the tag coordinate alone on that class. Since an $L$-bit tag provides at most $2^{L}$ outcomes, the clique size cannot exceed $2^{L}$.  $\blacksquare$
Corollary III.3footnotemark: footnotetext: : CIA3 (Counting converse).
@@
Theorem III.2 specialized to a single $K$-way ambiguity class gives $K\leq 2^{L}$, i.e., at least $\log_{2}K$ bits of side information.
Proposition III.4 (Equivalent formulations). The pair-injectivity obstruction of Theorem III.1 admits the following equivalent formulations on a finite source observed through a deterministic transcript with a finite auxiliary tag:
  • Global budget: $K\leq O\,T$ for latent size $K$, observation alphabet size $O$, and tag alphabet size TT.191919: OBS4
  • Fiber injectivity: $|\mathcal{F}_{y}|\leq|\mathcal{T}|$ for each observation fiber $\mathcal{F}_{y}=\{x:Y(x)=y\}$.202020: OBS3
  • Conditional entropy: $H(X\mid Y)=\log_{2}K$ on a $K$-way class with constant observation, and $H(T\mid Y)\geq\log_{2}K$.212121: PRB47, PRB55
  • Decoder output: $H(\hat{X})\leq H(Y,T)\leq H(X)$, with nonnegative gap to alphabet ceilings.222222: PRB73-74, PRB82-83, PRB90, PRB93-95
  • Mass-weighted clique: $\sum_{x\in S}p(x)\log_{2}\frac{1}{p(x)}\leq h_{2}(p_{S})+p_{S}\log_{2}|\mathcal{T}|$ for a success-set clique SS.232323: PRB55
  • Data processing: $H(\kappa(Y,T))\leq H(Y,T)$ for any deterministic coarsening $\kappa$.242424: PRB68, PRB81, PRB83
@@
Proof sketch. Each formulation follows from Theorem III.1 by restricting the injectivity condition to the relevant domain (fiber, clique, full source) and applying standard entropy identities. The confusability, counting, conditional-entropy, decoder-output, and finite-gap statements are therefore different normal forms of the same obstruction: a surviving $K$-way ambiguity class requires a budget of at least $\log_{2}K$ bits to be resolved exactly.252525: OBS1, OBS5, CIA3
III-A Finite-error extension
@@
The main body of the paper studies zero error. The same deterministic finite model also supports a budgeted finite-error extension, which should be read as a deterministic finite analogue of the classical Fano line of argument rather than as a separate asymptotic coding theorem [11, 12, 7, 13].
Proposition III.4footnotemark: footnotetext: : PRB85, PRB87 (Budgeted finite-error converse).
@@
IV Graph Characterization of Partial Views
@@
[…]
IV-A A Running Example: The Binary Square
@@
View 1 reveals $x_{1}$ and View 2 reveals $x_{2}$.
@@
Two states are confusable when at least one view fails to separate them. View 1 reveals $x_{1}$, so $(0,0)\sim(0,1)$ and $(1,0)\sim(1,1)$. View 2 reveals $x_{2}$, so $(0,0)\sim(1,0)$ and $(0,1)\sim(1,1)$. The confusability graph is therefore a 4-cycle:
@@
where opposite corners are not adjacent.
@@
The graph is 2-colorable (e.g., color by parity $c(x_{1},x_{2})=x_{1}\oplus x_{2}$), so exact recovery is possible with a 2-ary auxiliary tag. With one tag, the best exact success probability under the uniform source is $1/2$.
@@
Coding-theoretic interpretation. The parity coloring $c(x_{1},x_{2})=x_{1}\oplus x_{2}$ is a syndrome bit: with $T=2$ we decode perfectly, with $T=1$ we cannot distinguish adjacent vertices. The confusability graph is the confusion graph of a code with parity-check H=[1 1]H=[1\;1].
@@
Figure 1 contrasts this structured failure with clique-shaped failure. The theorems below generalize this to arbitrary multi-fact partial-view systems.
IV-B General Graph Characterization
This section turns the one-shot obstruction into the paper’s main structural object: the confusability graph induced by admissible partial views, where two tuples are confusable when allowed observations fail to separate them. For $d$ facts over alphabet size $q$, explicit graph materialization is exponentially large ($q^{d}$ vertices), so the formalization uses an implicit oracle interface.272727: MFT113 A naive explicit construction ranges over at most $q^{2d}$ ordered state pairs.282828: MFT116-117 The formalization packages adjacency as a Boolean confusability oracle and an equivalent agreement-set oracle, computable in $O(d+\sum_{\ell}|V_{\ell}|)$ time by computing the agreement set and scanning views.292929: MFT114-115, MFT132-135
@@ IV-B General Graph Characterization
The deterministic restriction is also not fundamental at the graph-construction level. The Lean development formalizes a support-overlap extension in which each location may emit any observation from a finite support set, and two states are confusable when some admissible location has overlapping observation support on those states.303030: MFT110-112
@@
The exact coordinate-subset model is not graph-universal. Equality at a view equals containment of that view in the coordinate-agreement set, so confusability depends only on which coordinates agree.313131: MFT118-119, MFT122 From any view family we obtain an upward-closed family $\mathcal{U}:=\{S\subseteq[F]:\exists\ell,V_{\ell}\subseteq S\}$ such that two tuples are adjacent iff their agreement set lies in $\mathcal{U}$. Conversely, every upward-closed family arises from some coordinate-view architecture.323232: MFT125-128 When the alphabet has at least two symbols, every proper coordinate subset occurs as the agreement set of some distinct pair of tuples.333333: MFT129-131 Coordinatewise alphabet permutations preserve the confusability relation and induce graph automorphisms.343434: MFT120-124 The model realizes vertex-transitive graphs on the full tuple space; non-vertex-transitive graphs such as a three-vertex path cannot arise as full confusability graphs. Classification up to unlabeled graph isomorphism remains open.
Theorem IV.1footnotemark: footnotetext: : MFT3-4, MFT20 (Exact recovery as graph colorability).
@@
For a finite multi-fact latent tuple observed through a finite family of partial-coordinate views, exact zero-error recovery with a $T$-ary auxiliary tag exists if and only if the induced confusability graph is $T$-colorable [14, 15].
@@
Necessity. Suppose exact recovery with $T$ tags is possible. If two latent tuples are adjacent in $G_{\mathrm{conf}}$, then by definition some allowed observation leaves them indistinguishable. Exact recovery therefore cannot assign them the same tag, because the joint observation-tag pair would then coincide on two distinct latent tuples and Theorem III.1 would be violated. Hence the tag assignment is a proper $T$-coloring of $G_{\mathrm{conf}}$.
Sufficiency. Suppose $c$ is a proper $T$-coloring of $G_{\mathrm{conf}}$. Define the auxiliary tag of a latent tuple to be its color. If two latent tuples produce the same observation and the same color, then they are confusable and adjacent, contradicting proper colorability. Therefore the joint observation-color map is injective, and Theorem III.1 yields an exact decoder.  $\blacksquare$
This statement is packaged through an explicit confusability graph object, and for $n$ repeated copies the full block confusability graph is identified with the $n$-fold strong power of the one-shot confusability graph.363636: MFT34-35
Theorem IV.2footnotemark: footnotetext: : MFT9 (Success-set characterization).
@@
Restrict the argument of Theorem IV.1 to the success set. Exactness is required only on $S$, so the coloring condition is required only on confusable pairs inside $S$.  $\blacksquare$
Theorem IV.3footnotemark: footnotetext: : MFT14-15 (Exact finite-error success budget).
@@
$M_{T}:=\max\{|S|:S\text{ induces a $T$-colorable subgraph}\}.$
@@
Theorem IV.2 identifies exact success sets with induced $T$-colorable subgraphs. Since the latent alphabet is finite, a maximum-cardinality such set exists. The mechanized characterization exposes this optimum directly. In the binary four-cycle witness, the optimum at $T=1$ is exactly two states, so the best uniform exact success probability is $2/4=1/2$.  $\blacksquare$
The same characterization is also available in weighted form. For any finite source weight $w$ on the latent alphabet, the maximum mass of a $T$-colorable induced subgraph is the exact finite success value under budget $T$. This is the weighted version of the same structural claim: under limited budget, the best exact decoder is the largest colorable safe subworld allowed by the induced failure graph.393939: MFT18-19
Equivalently, this optimum can be packaged as an explicit exact finite rate-distortion value for the extended model at tag budget $T$: every exact success set has weighted mass at most this value, and some exact decoder attains it. Thus the extended model no longer has only a converse inequality; at the one-shot level it has an exact weighted success law rather than merely an upper bound.404040: MFT23-24
Theorem IV.4footnotemark: footnotetext: : MFT5-8, MFT10 (First non-clique witness).
@@
Theorem IV.5footnotemark: footnotetext: : MFT11-13, MFT16 (Block-composition law).
@@
The strong product of graphs $G$ and $H$, denoted $G\boxtimes H$, is defined as follows. The vertex set is $V(G)\times V(H)$. Two vertices $(u,v)$ and $(u^{\prime},v^{\prime})$ are adjacent exactly when either $u=u^{\prime}$ and $v\sim v^{\prime}$, or $u\sim u^{\prime}$ and $v=v^{\prime}$, or $u\sim u^{\prime}$ and $v\sim v^{\prime}$. The block-composition theorem is model-specific because the admissible product views generate exactly these three adjacency cases [16, 14].
@@
The product confusability graph is exactly the strong product of the two component graphs.434343: MFT25-26 Clique-based lower bounds propagate to products: if the first component contains a confusability clique of size $c_{1}$ and the second contains one of size $c_{2}$, then the product contains a clique of size $c_{1}c_{2}$.444444: MFT17
The induced graph class has four structural features: it contains non-clique graphs (binary square witness), the cluster-collapse subclass on which the equality theorem is exact, closure under block-composition (strong product), and is exactly a monotone agreement-set class.
@@
Let $W$ be the binary square from Theorem IV.4. The $m$-fold product has confusability graph $C_{4}^{\boxtimes m}$ and exact tag budget $T_{\mathrm{exact}}(W^{\otimes m})=2^{m}$. This gives an infinite family of non-clique examples on which the graph object, exact budget, and strong-product law interact nontrivially.454545: MFT11-13, MFT16-17, MFT29-30
@@
V Block Composition, Strong Powers, and Asymptotic Capacity